OpenStudy (anonymous):

Solve the equation by factoring 2x-17=9/x

hero (hero):

2x^2-17x - 9 = 0

hero (hero):

2x^2-18x+x - 9 = 0

OpenStudy (anonymous):

ohhh ok

hero (hero):

2x(x-9)+1(x-9) = 0 (x-9)(2x+1) = 0

hero (hero):

x = 9,-1/2

OpenStudy (anonymous):

I see, thank you so much